OpenStudy (anonymous):

Find the values of x in the following equation: |-5x + -2| = 12

OpenStudy (anonymous):

x=2 ?

jhonyy9 (jhonyy9):

-5x-2=12 --- x=14/(-5) -5x-2=-12 --- x= -10/-5 = 2